Accelerator Update for June 21, 2006 - June 22, 2006
Wednesday June 21
The day shift began with Operations monitoring a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) in access, with the Antiproton Source (Pbar) not stacking due to Booster work in the 400MeV area, with the Main Injector (MI) and Recycler conducting studies, and with MiniBooNE and NuMI off due to Booster work.
Booster experts completed their work in the 400 MeV area. Operations established beam to the Booster dump at 11:40 AM.
By 12:15 PM, Operations had established beam to MiniBooNE and Pbar had resumed stacking.
Operations put the TeV and CDF power supplies into a dry squeeze at 3:52 PM.
At 4:53 PM, a Booster chopper (B:CHOP) tripped off. Experts accessed the Linac at 5:38 PM and continued their investigations.
Booster experts reported at 6:31 PM that they replaced a bad cable for the chopper.
Beam resumed at 6:37 PM, but it was erratic.
At 7:08 PM, Operations reported that the chopper tube wasn’t firing properly. Operations called in an expert who tightened all the connections. Beam resumed at 9:42 PM.
Operations began preparing the Recycler around 9:56 PM to stash antiprotons.
The D0 access ended at 11:23 PM.
Thursday June 22
The midnight shift began with Pbar and the Recycler stashing, with normal operations for MI, and with MiniBooNE and NuMI taking beam.
Operations put the TeV into a dry squeeze at 12:07 AM.
The Recycler stash ended at 12:32 AM. Pbar resumed stacking at 12:41 AM.
Operations put the TeV into a wet squeeze at 1:21 AM.
Operations put the TeV into shot setup at 2:15 AM.
At 4:35 AM, the TeV suffered a quench at sector F4. After a good deal of investigation, TeV experts felt that the quench was due to the orbit stabilization program.
The Plans for Thursday
The plans for today are to stack and stash so a store can get established. The Run Coordinator will allow TeV studies.
Comments and Suggestions
|last modified 06/22/2006 email Fermilab|